Reynoso Corporation manufactures titanium and aluminum tennis racquets. Reynoso’s total overhead costs consist of assembly costs and inspection costs. The following information is available:

Cost Titanium Aluminum Total Cost
Assembly 500 mach. hours 500 mach. hours $68000
Inspections 350 150 $106000
2100 labor hours 1900 labor hours



Reynoso is considering switching from one overhead rate based on labor hours to activity-based costing.
Using activity-based costing, how much assembly cost is assigned to titanium racquets?

$19950.

$34000.

$47600.

$35700.
Answer #1

Total machine hours=(500+500)=1000 machine hours

activity rate=Respective cost/Respective activity

=68000/1000

=$68 per machine hour

Hence assembly cost assigned to titanium racquets=(68*500)

=$34000